OpenStudy (anonymous):

log(5)-log(2x)=1...what is x?

OpenStudy (anonymous):

start with \[\log(\frac{5}{2x})=1\]

OpenStudy (anonymous):

then assuming this is log base ten rewrite as \[\frac{5}{2x}=10\] and solve for x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw